OpenStudy (anonymous):

Tyler works in retail where she earns $10.50 per hour. She also earns $15 for every new credit account she gets customers to open and she gets a weekly stipend of $50 for clothing. Tyler knows her weekly earnings can be shown with the following expression: 10.50h + 15c + 50 Part A: Identify a coefficient, a variable, and a constant in this expression. Part B: If Tyler works for 37 hours and opens 12 new credit accounts, how much does she earn? Show your work to receive full credit.

OpenStudy (anonymous):

Part A: A coefficient is a number attached to a variable, so for your equation 10.50h+15c+50 you could say 10.50 is a coefficient and h is the variable or you could say 15 is a coefficient and c is a variable. Either way 10.50 and 15 are the coefficients and h and c are the variables. 50 is not attached to any variable making it a constant. Part B: is a simple plug and chug! you have the equation 10.50h + 15c+ 50 h= number of hours worked c= number of credit accounts opened using this information you can plug in the 37 hours for h and 12 credits for c your equation should look like: 10.50(37) + 15(12)+ 50 = (388.5) + (180) +50 =$618.50 multiplying 10.50 by 37 will give you the 388.5 multiplying 15 by 12 will give you the 180 then you can add up all your totals and it will give you her earnings of $618.50